# MHONGOOSE — A MeerKAT Nearby Galaxy H i Survey

### v1

@@ 4 Observations
If J1939–6342 was used as the primary calibrator, two additional three-minute observations of a polarisation calibrator were inserted between different cycles. If J0408–6545 was used, a polarization calibrator was observed for three minutes at the end of every phase calibrator-target cycle. This increased number of polarization calibrator observations reflects the less well-characterised polarisation properties of J0408–6545. However, since we are only concerned with the Stokes $I$ data, we will not discuss the polarization aspects here.
@@ 5 Data Reduction
After an observation was completed and deposited in the SARAO MeerKAT archive, we extracted the data covering the channel range 16384 to 26383 (10,000 channels) from the 32k narrowband data set retaining only the HH and VV polarisations. This covers the topocentric frequency range 1390.0–1422.7 MHz. These data were binned by two channels leading to a 5000-channel measurement set with a channel width of 6.53 kHz (1.4 km s-1). The data were then processed using the CARACal data reduction pipeline [^4] CARACal and its related software packages can be found in the repository collection at https://github.com/caracal-pipeline/ (Józsa et al. 2020). As the observations cover multiple years, and MHONGOOSE observations took place simultaneously with CARACal development, several (pre-)releases of CARACal were used. Where necessary data were re-reduced to ensure consistent outcomes.
CARACal is a python-based, containerised pipeline that uses a “best-of” approach to link together reduction and analysis tasks and applications from various radio astronomy packages to optimally reduce the data, passing it seamlessly between the various reduction stages using Stimela (Makhathini 2018). These stages consist of:
@@
We will describe these steps in some more details below. Much of the reduction procedure was developed in collaboration with the MeerKAT Fornax Survey (Serra et al. 2023). The description given here should therefore be regarded as complementary to that given in Serra et al. (2023). Unless stated otherwise, all software tasks mentioned below in Sec. 5.15.7 were executed as part of the CARACal pipeline environment. All reduction was done on the dedicated MeerGas cluster at ASTRON consisting of four computing nodes with 128 cores and 1 Tb of memory each.
@@ 5.1 Cross calibration
For each observation the primary calibrator was used to derive the delay, gain and bandpass calibration using baselines longer than 150m. To improve the signal-to-noise, the bandpass was smoothed using a 9-channel box-car filter. Any gaps in the bandpass (e.g., due to the flagging of the Galactic emission) were interpolated. A second cycle of flagging and calibrating then refined these solutions, which were subsequently applied to the secondary calibrator. The latter was used to track the gain and phase variations over time. These time-dependent solutions were also derived twice with an intermediate flagging step. The final solutions were then applied to the target observations.
@@ 5.4 $u=0$ flagging
Figure 4: Overview of the $u=0$ flagging procedure. The top row shows a 100-channel binned empty map with horizontal stripes of one target scan of a NGC 1566 observation on the left, and on the right the corresponding FFT amplitude image which shows a prominent vertical RFI feature at $u=0$. The middle row shows the same plots, but after removal of the $u=0$ stripe using our flagging procedure. The bottom row shows two zeroth-moment H i maps of NGC 1566, both created using a simple $2\sigma$ intensity cut. The bottom-left panel shows the map with the $u=0$ RFI still included. The bottom-right panel shows the same map but with the $u=0$ RFI removed.
@@ 5.6 Standard resolutions
555(1) Label used to refer to this resolution. (2) Robustness weighting value. (3) Taper used. ”0” mean no taper. (4) Pixel size of the data cubes and moment maps. (5) Average major axis beam size. (6) Average minor axis beam size. (7) Average position angle of major axis of beam. (8) Noise per channel. (9) $1\sigma$, 1-channel column density sensitivity. (10) $3\sigma$, 16 km s-1 column density sensitivity. (11) H i mass sensitivity for a $3\sigma$, 50 km s-1 unresolved source at 10 Mpc.
@@ 5.7 Creating the full-depth cubes
For the full-depth observations, we combine the ten individual 5.5h observations and create data cubes measuring $1.5^{\circ}\times 1.5^{\circ}$ with a velocity range from $-500$ km s-1 to $+500$ km s-1 with respect to the central velocity of the target galaxy. We choose to image an area larger than the usual primary beam of $\sim 1^{\circ}$ due to the high sensitivity of the observations which enable H i detections at a large distance from the pointing center.
@@
Figure 6: Natural-weighted noise values measured over a 1.4 km s-1 channel as a function of integration time for the observations of ESO 300-G014 (J0309–41). The black, filled points indicate the measured noise level. The blue line shows the expected theoretical noise levels with a slope of $-0.5$. A linear fit to the observed noise levels gives a slope of $-0.504$. The red star indicates the target sensitivity from de Blok et al. (2016).
@@ 5.8 Creating full-depth moment maps
The second-moment map is commonly referred to as a velocity dispersion map. However, strictly speaking, second-moment values only represent a physically meaningful velocity dispersion in cases where the H i emission line profile is well approximated by a single Gaussian [^6] Note that here we adopt the informal definition that the second moment is equal to the dispersion. Mathematically, the second moment is defined as the variance.. In all other cases (non-Gaussian profiles or multiple separate components) they simply quantify the spread of H i velocities along each line of sight.
We used SoFiA-2 to find the signal in the cubes and create the masks. The signal-to-noise ratio S/N in the cubes is high, and we found that we only needed a limited number of spatial kernels to isolate the signal. In fact, using spatial kernels larger than 1similar-toabsent1$\sim$ 1∼ 12222 beam sizes had the effect of introducing additional noise into the moment maps at the edges of the disks and these larger kernels were therefore avoided. We used cubes that were not corrected for the primary beam. This correction was applied to the zeroth-moment maps afterwards.
@@ 5.9 Examples of full-depth moment maps
A full scientific analysis of the full-depth data will be presented in forthcoming papers. For this reason, and as data are still being collected, we restrict ourselves here to presenting a short overview of the full-depth data for two of the galaxies that were completed early in the survey. We refer to Healy et al. (2023) and Maccagni et al. (in prep.) for an in-depth discussion. The main purpose of this overview is to give an indication of the high quality of the data.
@@
We plot contours of multiples of this value, thus clearly capturing in a single figure the detailed, compact high-column density distribution at the highest resolutions, and then transitioning into the diffuse, extended distribution of the low-column density H i at the lowest resolutions. Note the extended and irregular distribution of H i in the outer parts of NGC 1566, as well as the clumpy outer H i in NGC 5068. At column densities of $\sim 10^{20}$ cm-2 (corresponding with the edge of the main star-forming disks) very little of this irregular outer H i would be detected and both galaxies would give the impression of regular, symmetric disks; cf. the observations of NGC 1566 in Elagali et al. (2019) which concentrated on the high-column density disk. As we will show below, the outer H i components contain only a small fraction of the total H i mass of the galaxies, but they dramatically change the apparent morphology. As mentioned, more extensive analyses of these and other sample galaxies will be presented in future papers, but Fig. 7 already shows the power of these deep observations.
Figure 7: Multi-resolution zeroth-moment maps of NGC 1566 (left panel) and NGC 5068 (right panel) based on the full-depth data. For each resolution, contours of $2^{n}$ times the S/N = 3 value are shown (as listed in the legend). The resolutions shown are r05_t00 (red), r10_t00 (green), r15_t00 (blue), r05_t60 (purple), r10_t90 (olive). For each resolution two contours are shown, i.e., $n=0,1$, except for the highest resolution r05_t00 (red contour), where $n=0,1,2,...$ The r05_t00 moment map is shown in the background as a false colour image, with the column density levels indicated by the colour bar on the right. The beam sizes (colored according to the respective resolution) are shown in the bottom-right corner.
Figure 8: Example single-track moment maps using the r05_t00 resolution for two MHONGOOSE galaxies. Top row: NGC 1566 (J0419–54); bottom row: NGC 5068 (J1318–21). From left to right: (i): Combined grz-colour image from DECaLS. (ii): primary-beam corrected zeroth-moment or integrated H i intensity map. Contours as indicated in the Figure. The lowest contour represents S/N = 3, with subsequent contour levels increasing by a factor of two. (iii): First-moment map or intensity-weighted velocity field. Red colours indicate the receding side, blue colours the approaching side. The central velocity (listed in Tab. 1) is indicated by the thick contour. Other contours are spaced by 10 or 20 km s-1, as indicated in the Figure. (iv): Second-moment map: the colours show the range from 0 (light-blue) to 30 (red) km s-1. The lowest contour shows the 12 km s-1 level, and subsequent contours are spaced by 12 km s-1. The 24 km s-1 contour is shown in black. For both the first- and second-moment maps, pixels corresponding to values below the S/N = 3 column density in the zeroth-moment map were blanked. See Appendix A for a more extensive description.
@@ 5.10 Single-track moment maps
At the time of writing observations are still ongoing, and full-depth observations are not yet available for the entire sample. However, every galaxy in the sample has by now been observed for at least one 5.5h track. The results presented in the remainder of this paper are based on the single track data, unless otherwise stated. To give an overview of the sample and pursue initial sample-wide science, we created single-track moment maps at two of the standard resolutions r05_t00 and r15_t00 (beam size of $\sim 12^{\prime\prime}$ and $\sim 30^{\prime\prime}$, respectively) in an identical manner as for the full-depth data. We have limited this to two resolutions as the differences between the standard resolutions are not as pronounced here due to the ten times shorter integration time. Average beam sizes for these single-track observations are identical to those listed in Table 5, while column density and mass limits are a factor $\sqrt{10}$ less (cf. Fig. 6). These moment maps give a comprehensive overview of the entire MHONGOOSE sample as observed by MeerKAT.
In Fig. 8 we show the r05_t00 single-track moment maps for NGC 1566 (J0419–54) and NGC 5068 (J1318–21) which were also discussed in Sect. 5.9. In the Figure, we show the Dark Energy Camera Legacy Survey (DECaLS) [^7] http://www.legacysurvey.org (Dey et al. 2019) optical image, the zeroth-moment (integrated H i intensity) map, the first-moment (intensity-weighted mean velocity field) [^8] The colour map used for the velocity field is based on the perception-based colour maps available as part of the CosmosCanvas project: https://ascl.net/code/v/3580 and finally, the second-moment map. For both the first- and second-moment maps, pixels corresponding to values below the S/N = 3 column density in the zeroth-moment map were blanked. A comparison between Figs. 7 and 8 for these galaxies clearly show the improvement in sensitivity that we can expect from the full survey.
@@ 6 Global H i Profiles
The global H i profile or integrated H i spectrum of a galaxy shows the spatially-integrated H i flux density as a function of velocity or frequency. It is often used in scaling relation studies, such as the Tully-Fisher relation (Tully & Fisher 1977) and also in studies involving the total H i masses of galaxies. Global H i profiles are often measured using single-dish telescopes. These measure the total power, and therefore the total H i mass. Global H i profiles measured with synthesis radio telescope arrays are less straight-forward to interpret. Depending on the baseline distribution of the array and the extent of the source, these may suffer from the so-called “missing spacing” problem, which is caused by a lack of short baselines which prevents measurements of flux on larger angular scales. For galaxies that have an angular size similar or larger than the scales probed by the shortest baselines of the array this can lead to a severe underestimate of the total H i mass. In these cases, a comparison with the single-dish H i global profile is often used to determine whether any H i has been missed in the synthesis observations. For MeerKAT, with a shortest baseline of 29m, the largest scale at which signal can be detected is $\sim 20^{\prime}$.
@@ 6.1 Comparing the single-track and full-depth profiles
A first step is to compare the H i global profiles of the single-track data and full-depth data which differ by a factor of ten in observing time. This obviously means an increase in signal-to-noise for the full-depth data, but an interesting question is whether this can be used to deduce the existence of low-column density H i.
@@
6.2 Comparison with single dish profiles
We have shown that the longer integration times of the full-depth data allow us to detect a small amount of additional H i. However, this internal comparison does not address the “missing spacing” issue. Is MeerKAT able to give us an accurate measurement of the total H i mass of our galaxies?
As the MHONGOOSE sample is ultimately selected from HIPASS, single-dish Parkes observations are available with a beam size of $15^{\prime}$ and flux measurements for all but one of the sample galaxies are given in the HIPASS Bright Galaxy Catalog ( BGC; Koribalski et al. 2004). The one galaxy that is not in the BGC (J0429–27) falls below that catalogue’s flux limit. Instead we use the flux density value as listed in the Meyer et al. (2004) HIPASS Catalogue. In addition, 15 of our galaxies have been observed with the GBT with a beam size of $9^{\prime}$ (Sorgho et al. 2018; Sardone et al. 2021).
We can therefore test whether the MHONGOOSE observations manage to recover the total flux. As the differences between the flux densities derived from the single-track and full-depth data are only small (as discussed above), we use the single-track flux densities here, as this maximises the number of galaxies where a comparison can be made.

### v2

@@ 4 Observations
If J1939–6342 was used as the primary calibrator, two additional three-minute observations of a polarisation calibrator were inserted between different cycles. If J0408–6545 was used, a polarization calibrator was observed for three minutes at the end of every phase calibrator-target cycle. This increased number of polarization calibrator observations reflects the less well-characterized polarisation properties of J0408–6545. However, since we are only concerned with the Stokes $I$ data, we do not discuss the polarization aspects here.
@@ 5 Data reduction
After an observation was completed and deposited in the SARAO MeerKAT archive, we extracted the data covering the channel range 16384 to 26383 (10,000 channels) from the 32k narrowband data set retaining only the HH and VV polarisations. This covers the topocentric frequency range 1390.0–1422.7 MHz. These data were binned by two channels leading to a 5000-channel measurement set with a channel width of 6.53 kHz (1.4 km s-1). The data were then processed using the CARACal data reduction pipeline [^4] CARACal and its related software packages can be found in the repository collection at https://github.com/caracal-pipeline/ (Józsa et al. 2020). As the observations cover multiple years, and MHONGOOSE observations took place simultaneously with CARACal development, several (pre)releases of CARACal were used. Where necessary data were rereduced to ensure consistent outcomes.
CARACal is a Python-based, containerized pipeline that uses a “best-of” approach to link together reduction and analysis tasks and applications from various radio astronomy packages to optimally reduce the data, passing it seamlessly between the various reduction stages using Stimela (Makhathini 2018). These stages consist of:
@@
We describe these steps in more detail below. Much of the reduction procedure was developed in collaboration with the MeerKAT Fornax Survey (Serra et al. 2023). The description given here should therefore be regarded as complementary to that given in Serra et al. (2023). Unless stated otherwise, all software tasks mentioned below in Sec. 5.15.7 were executed as part of the CARACal pipeline environment. All reduction was done on the dedicated MeerGas cluster at ASTRON consisting of four computing nodes with 128 cores and 1 Tb of memory each.
@@ 5.1 Cross calibration
For each observation the primary calibrator was used to derive the delay, gain and bandpass calibration using baselines longer than 150m. To improve the signal-to-noise ratio (S/N), the bandpass was smoothed using a 9-channel box-car filter. Any gaps in the bandpass (e.g., due to the flagging of the Galactic emission) were interpolated. A second cycle of flagging and calibrating then refined these solutions, which were subsequently applied to the secondary calibrator. The latter was used to track the gain and phase variations over time. These time-dependent solutions were also derived twice with an intermediate flagging step. The final solutions were then applied to the target observations.
@@ 5.4 $u=0$ flagging
Figure 4: Overview of the $u=0$ flagging procedure. The top row shows a 100-channel binned empty map with horizontal stripes of one target scan of a NGC 1566 observation on the left, and on the right the corresponding FFT amplitude image, which shows a prominent vertical RFI feature at $u=0$. The middle row shows the same plots, but after removal of the $u=0$ stripe using our flagging procedure. The bottom row shows two zeroth-moment H i maps of NGC 1566, both created using a simple $2\sigma$ intensity cut. The bottom-left panel shows the map with the $u=0$ RFI still included. The bottom-right panel shows the same map but with the $u=0$ RFI removed.
@@ 5.6 Standard resolutions
555(1) Label used to refer to this resolution. (2) Robustness weighting value. (3) Taper used. ”0” mean no taper. (4) Pixel size of the data cubes and moment maps. (5) Average major axis beam size. (6) Average minor axis beam size. (7) Average position angle of major axis of beam. (8) Noise per channel. (9) $1\sigma$, one-channel column density sensitivity. (10) $3\sigma$, 16 km s-1 column density sensitivity. (11) H i mass sensitivity for a $3\sigma$, 50 km s-1 unresolved source at 10 Mpc.
@@ 5.7 Creating the full-depth cubes
For the full-depth observations, we combine the ten individual 5.5h observations and create data cubes measuring $1.5^{\circ}\times 1.5^{\circ}$ with a velocity range from $-500$ km s-1 to $+500$ km s-1 with respect to the central velocity of the target galaxy. We choose to image an area larger than the usual primary beam of $\sim 1^{\circ}$ due to the high sensitivity of the observations, which enable H i detections at a large distance from the pointing center.
@@
Figure 6: Natural-weighted noise values measured over a 1.4 km s-1 channel as a function of integration time for the observations of ESO 300-G014 (J0309–41). The black filled points indicate the measured noise level. The blue line shows the expected theoretical noise levels with a slope of $-0.5$. A linear fit to the observed noise levels gives a slope of $-0.504$. The red star indicates the target sensitivity from de Blok et al. (2016).
@@ 5.8 Creating full-depth moment maps
The second-moment map is commonly referred to as a velocity dispersion map. However, strictly speaking, second-moment values only represent a physically meaningful velocity dispersion in cases where the H i emission line profile is well approximated by a single Gaussian [^6] We note that here we adopt the informal definition that the second moment is equal to the dispersion. Mathematically, the second moment is defined as the variance.. In all other cases (non-Gaussian profiles or multiple separate components) they simply quantify the spread of H i velocities along each line of sight.
We used SoFiA-2 to find the signal in the cubes and create the masks. The S/N in the cubes is high, and we found that we only needed a limited number of spatial kernels to isolate the signal. In fact, using spatial kernels larger than 1similar-toabsent1$\sim$ 1∼ 12222 beam sizes had the effect of introducing additional noise into the moment maps at the edges of the disks and these larger kernels were therefore avoided. We used cubes that were not corrected for the primary beam. This correction was applied to the zeroth-moment maps afterwards.
@@ 5.9 Examples of full-depth moment maps
A full scientific analysis of the full-depth data will be presented in forthcoming papers. For this reason, and as data are still being collected, we restrict ourselves here to presenting a short overview of the full-depth data for two of the galaxies that were completed early in the survey. We refer to Healy et al. (2024) and Maccagni et al. (2024) for in-depth discussions. The purpose of this overview is to give an indication of the high quality of the data.
@@
We plot contours of multiples of this value, thus clearly capturing in a single figure the detailed, compact high-column density distribution at the highest resolutions, and then transitioning into the diffuse, extended distribution of the low-column density H i at the lowest resolutions. Note the extended and irregular distribution of H i in the outer parts of NGC 1566, as well as the clumpy outer H i in NGC 5068. At column densities of $\sim 10^{20}$ cm-2 (corresponding with the edge of the main star-forming disks) very little of this irregular outer H i would be detected and both galaxies would give the impression of regular, symmetric disks; cf. the observations of NGC 1566 in Elagali et al. (2019) which concentrated on the high-column density disk. As we show below, the outer H i components contain only a small fraction of the total H i mass of the galaxies, but they dramatically change the apparent morphology. As mentioned, more extensive analyses of these and other sample galaxies will be presented in future papers, but Fig. 7 already shows the power of these deep observations.
Figure 7: Multi-resolution zeroth-moment maps of NGC 1566 (left panel) and NGC 5068 (right panel) based on the full-depth data. For each resolution, contours of $2^{n}$ times the S/N = 3 value are shown (as listed in the legend). The resolutions shown are r05_t00 (red), r10_t00 (green), r15_t00 (blue), r05_t60 (purple), r10_t90 (olive). For each resolution two contours are shown, i.e., $n=0,1$, except for the highest resolution r05_t00 (red contour), where $n=0,1,2,...$ The r05_t00 moment map is shown in the background as a false color image, with the column density levels indicated by the color bar on the right. The beam sizes (colored according to the respective resolution) are shown in the bottom-right corner.
Figure 8: Example single-track moment maps using the r05_t00 resolution for two MHONGOOSE galaxies. Top row: NGC 1566 (J0419–54); bottom row: NGC 5068 (J1318–21). From left to right: (i): Combined grz-color image from DECaLS. (ii): primary-beam corrected zeroth-moment or integrated H i intensity map. Contours as indicated in the figure. The lowest contour represents S/N = 3, with subsequent contour levels increasing by a factor of two. (iii): First-moment map or intensity-weighted velocity field. Red colors indicate the receding side, blue colors the approaching side. The central velocity (listed in Tab. 1) is indicated by the thick contour. Other contours are spaced by 10 or 20 km s-1, as indicated in the Figure. (iv): Second-moment map: colors show the range from 0 (light-blue) to 30 (red) km s-1. The lowest contour shows the 12 km s-1 level, and subsequent contours are spaced by 12 km s-1. The 24 km s-1 contour is shown in black. For both the first- and second-moment maps, pixels corresponding to values below the S/N = 3 column density in the zeroth-moment map were blanked. See Appendix A for a more extensive description.
@@ 5.10 Single-track moment maps
Observations are still ongoing at the time of writing, and full-depth observations are not yet available for the entire sample. However, every galaxy in the sample has by now been observed for at least one 5.5h track. The results presented in the remainder of this paper are based on the single track data, unless otherwise stated. To give an overview of the sample and pursue initial sample-wide science, we created single-track moment maps at two of the standard resolutions r05_t00 and r15_t00 (beam size of $\sim 12^{\prime\prime}$ and $\sim 30^{\prime\prime}$, respectively) in an identical manner as for the full-depth data. We have limited this to two resolutions as the differences between the standard resolutions are not as pronounced here due to the ten times shorter integration time. Average beam sizes for these single-track observations are identical to those listed in Table 5, while column density and mass limits are a factor $\sqrt{10}$ less (cf. Fig. 6). These moment maps give a comprehensive overview of the entire MHONGOOSE sample as observed by MeerKAT.
In Fig. 8 we show the r05_t00 single-track moment maps for NGC 1566 (J0419–54) and NGC 5068 (J1318–21) which were also discussed in Sect. 5.9. In the Figure, we show the Dark Energy Camera Legacy Survey (DECaLS) [^7] http://www.legacysurvey.org (Dey et al. 2019) optical image, the zeroth-moment (integrated H i intensity) map, the first-moment (intensity-weighted mean velocity field) [^8] The color map used for the velocity field is based on the perception-based color maps available as part of the CosmosCanvas project: https://ascl.net/code/v/3580 and finally, the second-moment map. For both the first- and second-moment maps, pixels corresponding to values below the S/N = 3 column density in the zeroth-moment map were blanked. A comparison between Figs. 7 and 8 for these galaxies clearly show the improvement in sensitivity that we can expect from the full survey.
@@ 6 Global H i profiles
The global H i profile or integrated H i spectrum of a galaxy shows the spatially integrated H i flux density as a function of velocity or frequency. It is often used in scaling relation studies, such as the Tully-Fisher relation (Tully & Fisher 1977) and also in studies involving the total H i masses of galaxies. Global H i profiles are often measured using single-dish telescopes. These measure the total power, and therefore the total H i mass. Global H i profiles measured with synthesis radio telescope arrays are less straight-forward to interpret. Depending on the baseline distribution of the array and the extent of the source, these may suffer from the so-called “missing spacing” problem, which is caused by a lack of short baselines which prevents measurements of flux on larger angular scales. For galaxies that have an angular size similar or larger than the scales probed by the shortest baselines of the array this can lead to a severe underestimate of the total H i mass. In these cases, a comparison with the single-dish H i global profile is often used to determine whether any H i has been missed in the synthesis observations. For MeerKAT, with a shortest baseline of 29m, the largest scale at which signal can be detected is $\sim 20^{\prime}$.
@@ 6.1 Comparing the single-track and full-depth profiles
A first step is to compare the H i global profiles of the single-track data and full-depth data which differ by a factor of ten in observing time. This obviously means an increase in S/N for the full-depth data, but an interesting question is whether this can be used to deduce the existence of low-column density H i.
@@
6.2 Comparison with single-dish profiles
Above, we show that the longer integration times of the full-depth data allow us to detect a small amount of additional H i. However, this internal comparison does not address the “missing spacing” issue. A further question is therefore whether MeerKAT is able to give us an accurate measurement of the total H i mass of our galaxies.
As the MHONGOOSE sample is ultimately selected from HIPASS, single-dish Parkes observations are available with a beam size of $15^{\prime}$ and flux measurements for all but one of the sample galaxies are given in the HIPASS Bright Galaxy Catalog ( BGC; Koribalski et al. 2004). The one galaxy that is not in the BGC (J0429–27) falls below the flux limit of that catalog. Instead we use the flux density value as listed in the Meyer et al. (2004) HIPASS Catalogue. In addition, 15 of our galaxies have been observed with the GBT with a beam size of $9^{\prime}$ (Sorgho et al. 2018; Sardone et al. 2021).
We can therefore test whether the MHONGOOSE observations manage to recover the total flux. As the differences between the flux densities derived from the single-track and full-depth data are only small (as discussed above), we use the single-track flux densities here, as this maximizes the number of galaxies where a comparison can be made.